A person decides to use his bath tube water to generate electric ppwer to run a 40w bulb.the bath tube is located at a height of 10m from the ground and it hold 200litres of water.he installs a water driven wheel generator on the ground.at what rate should the water drain from the bath tube to light the bulb?hiw long can he keep the bulb on,if the bath tube was full initially?the efficiency of the generator is 90%.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Ankit
 
P = mgh/t
0.9.* mgh/t = 40
 
t = 0.9 * 200*9.8*10/40 = 441 sec
 
Rate = 200/441 = 0.453 lit/ sec
 
 
Regards
Arun(askIITiams forum expert)